OpenStudy (thebeatles1200):

Please someone help me! 10 5/12 - 3 9/10. What am I doing wrong? I keep getting 6 1/30 when its suppose to be 6 31/60.

OpenStudy (mathmale):

Showing us all your work would help us help you.

OpenStudy (thebeatles1200):

@mathmale Here is my work: (10-3)+(5/12-9/10) LCD=2x2x3x5 LCD=4x15=60 50/60-108/60 7+(50/60-108/60) since you cant subtract 50/60 from 108/60 properly I have to borrow from the whole 7 6(1 50/60-108/60) 1x60+50 60+50=110 6+110/60-108/60 6+2/60 6 1/30

OpenStudy (thebeatles1200):

@mathmale Please come back! I typed that all out for you :(

OpenStudy (thebeatles1200):

Ive spent way too long on this problem

OpenStudy (thatonegirl_):

Your LCD is 60 correct?

OpenStudy (thebeatles1200):

Yes

OpenStudy (thebeatles1200):

Isnt that right?

OpenStudy (thatonegirl_):

Yes. So you have to get a common denominator of 60 for both fractions, right?

OpenStudy (thatonegirl_):

\[\frac{ 5 }{ 12 }-\frac{ 9 }{ 10 }\] is this right?

OpenStudy (thebeatles1200):

Yes and then I multiply 5 by 10 and 9 by 12

OpenStudy (thebeatles1200):

Yes

OpenStudy (thebeatles1200):

Thats the original fractions

OpenStudy (thatonegirl_):

Oh so your denominator is going to be 120 then?

OpenStudy (thebeatles1200):

No...

OpenStudy (thebeatles1200):

I thought I only needed to multiply the top part of the fraction by the opposite bottoms

OpenStudy (thatonegirl_):

If you multiply the first fraction by 10 (both numerator and denominator), your denominator will be 120.

OpenStudy (thebeatles1200):

I tried doing that previously but my answer still came out to 6 1/30

OpenStudy (thatonegirl_):

When getting common denominators, you need to multiply top and bottom by the same thing or else you are changing the value of the fraction.

OpenStudy (thatonegirl_):

Okay well let's try it with 120 first, then we can do 60 later. Should be same answers.

OpenStudy (thatonegirl_):

\[\frac{ 5 }{ 12 }\times \frac{ 10 }{ 10 }\]

OpenStudy (thebeatles1200):

Okay so we're not finding LCD ?

OpenStudy (thebeatles1200):

Its 9/10 not 10/10

OpenStudy (thatonegirl_):

Well you already found the LCD, 60.

OpenStudy (thatonegirl_):

Yes but I'm multiplying by 10/10 (which is the same thing as 1) to change the denominator to 120.

OpenStudy (thebeatles1200):

Okay so we have the LCD now we multiply 5/12 by 10 to get 50/120 and 9/10 by 12 to get 108/120?

OpenStudy (thebeatles1200):

What????

OpenStudy (thatonegirl_):

Okay hang on a minute. So you have your LCD, 60. But where are you using that in your work?

OpenStudy (thebeatles1200):

Hold on Ill send a pic

OpenStudy (thebeatles1200):

Can you see it?

OpenStudy (thatonegirl_):

yup

OpenStudy (thebeatles1200):

Im not understanding what Im doing :(

OpenStudy (thebeatles1200):

Can someone go through the correct steps with me?

OpenStudy (thatonegirl_):

Okay I'll start from the beginning. You're getting confused on converting to the common denominator which is an easy thing to clear up :)

OpenStudy (mathmale):

Actually, you identified the LCD correctly minutes ago. You are subtracting 9/10 from 5/12. Can you convert these fractions so that both have the same denominator (60)? \[\frac{ 9 }{ 10 }\frac{ 6 }{ 6 }=?\]

OpenStudy (thebeatles1200):

Okay so I have it written as (10-3)+(5/12-9/10) and Ive found the LCD=60, now what do I do?

OpenStudy (mathmale):

Now you please change -9/10 so this fraction has a denom. of 60.

OpenStudy (thebeatles1200):

Wait, where did the 6 come from though?

OpenStudy (thatonegirl_):

\[\frac{ 5 }{ 12 }-\frac{ 9 }{ 10 }\] This is your original problem. (you already subtracted the whole numbers so we don't need to worry about that part.)

OpenStudy (thatonegirl_):

In order to combine fractions, they need the same bottom number (common denominator).

OpenStudy (thebeatles1200):

Yes which is the LCD, 60 right?

OpenStudy (mathmale):

If you are able to undrstand WHY the LCD is 60, then you can see that mult. 12 in the denom. of the first fraction by 6 will give you 60. Mult the numerator by 6 also. What do you get?

OpenStudy (thatonegirl_):

Yes perfect! You can really pick any number you want that both numbers are a factor of (so you could do 120 for the bottom number because 12 can go into 120 and 10 can also go into 120.) But the LCD is the easiest to work with, 60.

OpenStudy (thebeatles1200):

Okay so do I have 5/60-9/60 or 50/60-108/60

OpenStudy (mathmale):

(10-3)+(5/12-9/10) is fine, great, perfect. 10-3=7, and (separately), Combine 5/12 and -9/10. As before, \[\frac{ 5 }{ 12}\frac{ 5 }{ 5}-\frac{ 9 }{ 10 }\frac{ ? }{ ? }\]

OpenStudy (thebeatles1200):

Im sorry If Im giving you guys a hard time :/ Im so confused

OpenStudy (thatonegirl_):

Well in order to get 60 on the denominator what do you need to multiply the first fraction's denominator by? 12x___=60

OpenStudy (mathmale):

You must multiply 10 by what number to get 60...? Replace the ? marks with the proper numeral:\[\frac{ 5 }{ 12}\frac{ 5 }{ 5}-\frac{ 9 }{ 10 }\frac{ ? }{ ? }\]

OpenStudy (thatonegirl_):

No problem! It's understandable to be confused. You'll get it :)

OpenStudy (thebeatles1200):

so 5/12 times 5/5 and 9/10 times 9/9?

OpenStudy (thebeatles1200):

and 12x_=60 is 5

OpenStudy (mathmale):

your lcd is 60, so you would not want to multiply that denominator 10 by 9, would you? That'd give yu 90; you want 60. Try again.\[\frac{ 5 }{ 12}\frac{ 5 }{ 5}-\frac{ 9 }{ 10 }\frac{ ? }{ ? }\]

OpenStudy (thatonegirl_):

The first one is correct. To get the denominator to equal 60 you need to multiply 12 by 5. But what you do to the bottom, you must do to the top (very important rule!). So you multiply the numerator (top number) by 5 also. So what is the first "new" fraction equal to?

OpenStudy (thebeatles1200):

Ohh okay I see what your saying, I was too hung up on the 5 and 9 on top of the fraction its would be 9/10 time 6/6 then

OpenStudy (thatonegirl_):

Yup!

OpenStudy (thebeatles1200):

25/60 right?

OpenStudy (thatonegirl_):

Exactly!! How about the 2nd fraction?

OpenStudy (mathmale):

Yes. So, your 9/10 would become 54/60. So now you have 25/60 - 54.59, Combine these into 1 fraction.

OpenStudy (thebeatles1200):

Yes 54/60

OpenStudy (thebeatles1200):

Wait I cant do 25/60 - 54/60 though

OpenStudy (thebeatles1200):

I thought I have to borrow a 1 from the 7 whole number

OpenStudy (mathmale):

Yes. Glad you found my error! 25/60-54/60 = what? Your answer MUST have the LCD 60. What is 25-54?

OpenStudy (thebeatles1200):

so it would be 6+(1 25/60-54/60)

OpenStudy (mathmale):

You combined the integers first, and got 7; now you are finishing up combining the fractions, and have (25-54)/60. Simplify that. I think you're beginning teh borrowing process too soon.

OpenStudy (thebeatles1200):

So I would do 1 time 25 plus 60 to get 85/60 right?

OpenStudy (mathmale):

Let's slow down just a little bit. Please, combine 25/60 - 54/60.

OpenStudy (thebeatles1200):

then do 85/60-54/60 to get 6 31/60

OpenStudy (mathmale):

Please slow down. Too early to be borrowing!

OpenStudy (mathmale):

combine: 25/60 - 54/60.

OpenStudy (thebeatles1200):

Oh sorry, my teacher was having us doing borrowing in this lesson but I suppose you could subtract that to get a negetaive though

OpenStudy (mathmale):

Too early. please combine the fractions.

OpenStudy (thebeatles1200):

itd be -29/60

OpenStudy (thatonegirl_):

Actually you did get the right answer lol

OpenStudy (mathmale):

Yes, that's right. Now you have 7 - 29/60. Now you must borrow a 1, or 60/60, from that 7. Do that now.

OpenStudy (thebeatles1200):

I know :) Thanks!

OpenStudy (thatonegirl_):

So you can borrow then if that makes it easier for you :P

OpenStudy (thebeatles1200):

Thank you both for the help!

OpenStudy (thatonegirl_):

No problem, hope it helped some!

OpenStudy (mathmale):

Hope you'll review this problem solution so that you'll be better able to tackle the next similar problem.

OpenStudy (thebeatles1200):

It did tremendously! My teacher had shown us in class but I was looking at it the wrong way, thanks so much for going through it with me!

OpenStudy (thatonegirl_):

Ah no problem :) Good luck with the rest of your work!